Current and temperature measurements
I measured the following currents in the T90 (without film, connected to the laboratory power supply; in brackets the reference values according to the SPT Journal):
- Main switch OFF: 7,8 uA (7,0 uA)
- Main switch (LCD) ON: 44,5 uA (430 uA)
- Trigger switch (lightmeter, LEDs and LCD illumination) ON: 109 mA (190 mA)
- Serial shooting in motor mode HIGH: 600 mA (650 - 700 mA during film advance after exposure)
Additionally, I measured the surface temperature of the connectors on the DC/DC converter board with an IR thermometer.
The temperature remained constant even during serial operation.
I conclude that the problem was with the DC/DC converter and that replacing the board resolved it.
Possibilities to replace the viewfinder LEDs
A digit in the viewfinder display (a bar on one of the seven-segment LEDs) has failed (see the thread). This means that the displayed values can still be read.
I've looked into possible replacement options.
Position of the board with the LEDs.
Unit diassembled (on a spare parts mirror box).
The LED unit is attached to the FD board, which is soldered to the mainboard (blue arrow). It can be removed.
Alternatively, only the LED array could be separated from its circuit board. This can be done using hot air.
However, soldering it back on is a delicate process; the connections would have to be tinned and then re-soldered with hot air, out of sight.
Conclusion
- Both options are risky.
- Replacing the FD board with the LED array also involves replacing an IC that presumably controls the LEDs. I don't know the board versions or whether they're compatible, so it would be an experiment.
- Soldering and re-soldering the LED array is tricky, and if it's not done correctly, it could render the connection unusable.
- I'll therefore leave everything as it is. The failure of one digit doesn't justify the risk.
- It remains to be seen whether the failure is related to the previously defective DC/DC converter (possible overload). I'll continue to monitor this.